Solve quadratic equations using the squaring method.
x^2+8x-2=0????

Found 2 solutions by hokies, dabanfield:
Answer by hokies(65)

x = 4 + 3 * sqrt(2)
x = 4 - 3 * sqrt(2)

Answer by dabanfield(803)

Solve quadratic equations using the squaring method.
x^2+8x-2=0????
We need a perfect square, like (x+a)^2, on the left side of the equation. To determine what to add to each side to get a perfect square on the left, we look at the coefficient 8 on the x term. Half of 8 is 4. Squaring 4 we get 16 so we need to add 18 to both sides to get 16 on the left:
x^2 + 8x - 2 + 18 = 18
x^2 + 8x + 16 = 18
(x+4)*(x+4) = 18
(x+4)^2 = 18
Take the square root of both sides and we have:
x+4 = sqrt(18)
x = sqrt(18) - 4
x = 3*sqrt(2) - 2